### v1

# RecConv: Efficient Recursive Convolutions for Multi-Frequency Representations

Abstract
Recent advances in vision transformers (ViTs) have demonstrated the advantage of global modeling capabilities, prompting widespread integration of large-kernel convolutions for enlarging the effective receptive field (ERF). However, the quadratic scaling of parameter count and computational complexity (FLOPs) with respect to kernel size poses significant efficiency and optimization challenges. This paper introduces RecConv, a recursive decomposition strategy that efficiently constructs multi-frequency representations using small-kernel convolutions. RecConv establishes a linear relationship between parameter growth and decomposing levels which determines the effective kernel size $k\times 2^{\ell}$ for a base kernel $k$ and $\ell$ levels of decomposition, while maintaining constant FLOPs regardless of the ERF expansion. Specifically, RecConv achieves a parameter expansion of only $\ell+2$ times and a maximum FLOPs increase of $5/3$ times, compared to the exponential growth ($4^{\ell}$) of standard and depthwise convolutions. RecNeXt-M3 outperforms RepViT-M1.1 by 1.9 $AP^{box}$ on COCO with similar FLOPs. This innovation provides a promising avenue towards designing efficient and compact networks across various modalities. Codes and models can be found at https://github.com/suous/RecNeXt.
1 Introduction
[…]
Figure 1: Effective Receptive Field (ERF) and Model Performance. (a) Comparison of effective kernel sizes and performance metrics among RecNeXt and prominent lightweight models. Top-1 accuracy is evaluated on ImageNet-1K and the latency is assessed using an iPhone 13 running iOS 18, where all models are formatted into mlmodel. $\dagger$” denotes pretraining at $256\times 256$ resolution and $\ast$” indicates pretraining with hard knowledge distillation. (b) The effective receptive fields of RepViT [64], FastViT [60], RepNeXt [72], and RecNeXt, where all models are pretrained with hard knowledge distillation. RecNeXt exhibits the broadest ERF, maintaining competitive latency and superior performance without reliance on structural reparameterization (SRP).
Convolutional Neural Networks (CNNs) [28, 27, 18, 49] have been foundational in computer vision for years, exploiting their intrinsic locality and translation equivariance [12]. Nonetheless, the recent emergence of vision transformers (ViTs) has demonstrated that establishing global context through self-attention mechanisms can lead to superior performance on various computer vision benchmarks [12, 2, 73]. Following the success of ViTs, extensive research has validated the effectiveness of large kernel convolutions in augmenting model capabilities through expanded receptive fields. ConvNeXt [38] and RepLKNet [9] achieved performance comparable or superior to Swin Transformers [37] when aligned with contemporary design paradigms. SLaK [36] and PeLK [3] explored kernels beyond $51\times 51$ and $101\times 101$ through dynamic sparsity and parameter sharing. Moreover, large-kernel CNNs have shown superiority as teachers in knowledge distillation for compact CNNs [24].
Although large-kernel convolutions provide superior performance, the quadratic scaling of parameter count and computational complexity relative to kernel size presents substantial efficiency challenges and optimization difficulties, restricting their practical deployment on resource-constrained devices. Consequently, small-kernel convolutions are still preferred [64, 65, 51, 34] in such scenarios considering their optimized algorithms and compact weights. Recent studies have explored large-kernel convolutions in resource-constrained environments. FastViT [60], StarNet [39], and EfficientMod [40] employed $7\times 7$ depthwise convolutions to capture long-range dependencies. EdgeNeXt [41] introduced adaptive kernel mechanisms for multi-level feature extraction. MobileMamba [17] enhanced multi-scale representations through a multi-receptive field feature interaction module. RepNeXt [72] employed multi-branch efficient operations to replace the $11\times 1$ 11 depthwise convolution. However, the application of even larger convolution kernels in resource-constrained settings remains unexplored.
OctConv [5] leveraged octave convolutions to enable multi-frequency feature representation. WTConv [13] incorporated small-kernel convolutions with cascade wavelet decomposition to obtain global receptive fields. While effective, their implementation requires substantial amount of parameters and resource-intensive operations. Motivated by the multi-frequency response observed in OctConv and WTConv, we propose RecConv — a recursively decomposed multi-frequency convolution optimized for resource-constrained mobile vision applications. Based on RecConv, we introduce RecNeXt, a simple yet effective vision backbone that delivers competitive efficiency with superior performance. Our contributions can be concluded as follows:
  • Inspired by WTConv [13], we develop a recursive decomposition strategy to generate multi-frequency representations. This approach establishes a linear relationship between parameter growth and decomposing levels, while maintaining constant FLOPs, making it highly adaptable for designing efficient networks with large receptive fields across diverse modalities.
  • Following this design principle, we construct RecConv as a plug-and-play module that can be seamlessly integrated into existing computer vision architectures.
  • Leveraging RecConv, we introduce RecNeXt, an efficient and effective vision backbone equipped with large receptive fields and optimized towards resource-constrained mobile vision applications.
  • Extensive experiments across various vision benchmarks demonstrate the superiority of RecNeXt over current leading approaches without relying on neural architecture search (NAS) or stractural reparameterization (SRP).
2 Related Work
Large Kernel CNNs: Large kernels have been explored in early CNNs [27, 53], but fell out of favor after VGG [52]. ConvNeXt [38] and ConvMixer [59] embraced contemporary design philosophy with large-kernel depthwise convolutions, replicating the global perspective of ViTs [12] and MLP-Mixers [58]. MogaNet [30] incorporated multi-scale spatial aggregation blocks with dilated convolutions to capture discriminative features. SKNet [31] and LSKNet [33] employed multi-branch selective mechanisms along the channel or spatial dimension. RepLKNet [9] achieved performance comparable to Swin Transformers [37] by expanding kernel size to $31\times 31$ with SRP. UniRepLKNet [10] demonstrated the universal applicability of large-kernel designs across various modalities. SLaK [36] improved performance by expanding kernels beyond $51\times 51$ using stripe convolutions and dynamic sparsity. PeLK [3] explored kernels up to $101\times 101$ in a human-like pattern without performance saturation. GFNet [48] leveraged Fourier transforms to capture long-term spatial dependencies in the frequency domain. WTConv [13] replaced large-kernel convolutions with a set of small-kernel convolutions through wavelet decomposition. Furthermore, $n\times n$ convolutions are often factorized into sequential or parallel $1\times n$ and $n\times 1$ convolutions for efficiency [54, 55, 46, 15, 69]. Additionally, LargeKernel3D [7] and ModernTCN [11] extended large kernel design to 3D networks and time series analysis.
Efficient Networks: Developing efficient networks for resource-constrained vision applications has attracted significant attention [22, 49, 65]. MobileNets [22, 50, 21] proposed depthwise separable convolutions and inverted residual blocks to optimize the efficiency-accuracy trade-off. GhostNet [16], FasterNet [4], and SHViT [70] employed partial channel operations for parameter reduction and computational efficiency. MobileOne [61], FastViT [60], and RepViT [64] utilized structural reparameterization (SRP) to improve feature diversity without increasing the inference latency. MnasNet [57] and EfficientNet [56] leveraged neural architecture search (NAS) to automatically discover efficient architectures. MobileViTs [42, 43, 63], EdgeViTs [45], and EfficientFormers [32, 34] designed hybrid architectures to balance performance and efficiency. Mobile-Former [6] and ParC-Net [71] introduced multi-branch feature integration to improve representational capacity. SwiftFormer [51] and LightViT [23] constructed linear attention mechanisms to replace quadratic matrix multiplications. FastViT [60], StarNet [39], and EfficientMod [40] adopted $7\times 7$ depthwise convolutions to capture long-range dependencies. EdgeNeXt [41] and RepNeXt [72] introduced multi-branch efficient operations to obtain large-kernel receptive fields. MobileMamba [17] incorporated long-range wavelet transform-enhanced mamba and efficient multi-kernel depthwise convolution to enhance multi-scale perception.
The major perspective of our work is inspired by WTConv [13], while our design focuses on resource-constrained applications. And there are three major differences between WTConv and our proposed method: (1) We adopt a shared and learnable downsampling layer using a depthwise convolution with a stride of 2. (2) We maintain the same channel dimension across different scales for structural consistency and computational efficiency. (3) We use bilinear interpolation as an efficient upsampling operation to reduce parameter count and memory footprint.
3 Method
3.1 Overall Architecture
Table 1: Classification performance on ImgeNet-1K. Following [32, 64], latency is measured on an iPhone 13 with models compiled by Core ML Tools. Similar to [14], throughput is tested on an Nvidia RTX3090 GPU with maximum power-of-two batch size that fits in memory. $\dagger$” denotes the evaluation image size is 256.
Model Type Params MACs Latency $\downarrow$ Throughput $\uparrow$ Top1 $\uparrow$ Auxiliary
(M) (G) (ms) (im/s) (%) NAS SRP
MobileViG-Ti [44] CNN-GNN 5.2 0.7 1.4 4337 75.7
SwiftFormer-XS [51] Hybrid 3.5 0.6 1.5 4304 75.7
EfficientFormerV2-S0 [34] Hybrid 3.5 0.4 1.4 1274 75.7
FastViT-T8$\dagger$ [60] Hybrid 3.6 0.7 1.3 3909 76.7
RepViT-M0.9 [64] CONV 5.1 0.8 1.3 4817 78.7
RepNeXt-M1 [72] CONV 4.8 0.8 1.3 3885 78.8
EfficientFormerV2-S1 [34] Hybrid 6.1 0.7 1.6 1153 79.0
RepViT-M1.0 [64] CONV 6.8 1.1 1.4 3910 80.0
RepNeXt-M2 [72] CONV 6.5 1.1 1.4 3198 80.1
RecNeXt-M1 CONV 5.2 0.9 1.4 384 79.2
RecNeXt-M2 CONV 6.8 1.2 1.5 325 80.3
MobileViG-S [44] CNN-GNN 7.2 1.0 1.6 2985 78.2
SwiftFormer-S [51] Hybrid 6.1 1.0 1.8 3376 78.5
EfficientFormer-L1 [32] Hybrid 12.3 1.3 1.8 3360 79.2
FastViT-T12$\dagger$ [60] Hybrid 6.8 1.4 1.8 3182 80.3
RepViT-M1.1 [64] CONV 8.2 1.3 1.5 3604 80.7
RepNeXt-M3 [72] CONV 7.8 1.3 1.5 2903 80.7
RecNeXt-M3 CONV 8.2 1.4 1.6 314 80.9
MobileViG-M [44] CNN-GNN 14.0 1.5 2.0 2491 80.6
FastViT-S12$\dagger$ [60] Hybrid 8.8 1.8 2.0 2313 80.9
SwiftFormer-L1 [51] Hybrid 12.1 1.6 2.2 2576 80.9
EfficientFormerV2-S2 [34] Hybrid 12.6 1.3 2.3 611 81.6
FastViT-SA12$\dagger$ [60] Hybrid 10.9 1.9 2.3 2181 81.9
RepViT-M1.5 [64] CONV 14.0 2.3 1.9 2151 82.3
RepNeXt-M4 [72] CONV 13.3 2.3 1.9 1745 82.3
RecNeXt-M4 CONV 14.1 2.4 2.4 169 82.5
EfficientFormer-L3 [32] Hybrid 31.3 3.9 3.4 1422 82.4
MobileViG-B [44] CNN-GNN 26.7 2.8 3.5 1446 82.6
SwiftFormer-L3 [51] Hybrid 28.5 4.0 3.4 1474 83.0
EfficientFormer-L7 [32] Hybrid 82.1 10.2 7.1 619 83.3
EfficientFormerV2-L [34] Hybrid 26.1 2.6 3.8 399 83.3
RepViT-M2.3 [64] CONV 22.9 4.5 2.8 1184 83.3
FastViT-SA24$\dagger$ [60] Hybrid 20.6 3.8 3.4 1128 83.4
RepNeXt-M5 [72] CONV 21.7 4.5 2.8 978 83.3
RecNeXt-M5 CONV 22.9 4.7 3.4 104 83.3
Figure 2: Recursive Convolution Module (RecConv). The left side illustrates a MetaNeXt block abstracted from previous works [68, 38, 69]. The right side highlights the RecConv details, where feature maps are recursively decomposed into multiple spatial scales using a shared depthwise convolutional downsampling layer. Each scale applies a small-kernel depthwise convolution to generate features of varying spatial sizes, which are then upsampled and aggregated, collectively replicating the large-kernel receptive field while preserving computational and parameter efficiency. This design strategy accommodates various functions and can be extended to other modalities.
The RecNeXt architecture builds upon RepViT [64] and RepNeXt [72], adhering to the four-stage framework of conventional CNNs [18] and hierarchical ViTs [37]. Starting with two $3\times 3$ convolutions with a stride of $2$ [64, 34, 67], progressively enhancing semantic representations while reducing spatial dimensions. The micro-blocks adopt the MetaNeXt design [38, 69], incorporating a token mixer for spatial feature extraction, a channel mixer for visual semantic interaction, a normalization layer [25] for training stability, and a shortcut connection [18] to facilitate smoother optimization [29].
$Y=X+\mathrm{ChannelMixer}\big{(}\mathrm{Norm}(\mathrm{TokenMixer}(X))\big{)},$ (1)
where $X,Y\in\mathbb{R}^{B\times C\times H\times W}$ with $B$ represents batch size, $C$ denotes channel number, and $H$ and $W$ indicate image height and width, respectively. $\mathrm{Norm}(\cdot)$ denotes the Batch Normalization (BN) layer [25], which can be fused into the previous convolution layer. $\mathrm{TokenMixer}(\cdot)$ implements the RecConv and $\mathrm{ChannelMixer}(X)=\mathrm{Conv_{1\times 1,\downarrow}}\big{(}\sigma(\mathrm{Conv_{1\times 1,\uparrow}}(X))\big{)}$ is a channel MLP module comprising two fully-connected layers with a GELU [20] activate function $\sigma$ in between, resembling the feed-forward network in a Transformer [62]. $\mathrm{Conv_{1\times 1,\uparrow}}$ and $\mathrm{Conv_{1\times 1,\downarrow}}$ stand for $1\times 1$ pointwise convolutions for expanding and squeezing feature maps, respectively.
The downsampling layer between each stage is a modified version of the MetaNeXt block [38, 69], where the shortcut connection bypasses the channel mixer.
$\begin{split}\hat{X}&=\mathrm{Norm}(\mathrm{TokenMixer}(X)),\\ Y&=\hat{X}+\mathrm{ChannelMixer}(\hat{X}),\end{split}$ (2)
where $\hat{X},Y\in\mathbb{R}^{B\times 2C\times H/2\times W/2}$. $\mathrm{TokenMixer}(\cdot)$ is a $7\times 7$ depthwise convolution with stride $2$ for downsampling.
3.2 Recursive Convolution
The right side of Figure 2 spotlights the RecConv details, where initial feature maps are recursively decomposed into multiple spatial scales via a shared depthwise convolutional downsampling layer. Each scale employs a small-kernel depthwise convolution to extract features at varying spatial resolutions, which are then progressively fused back into the original feature map using bilinear upsampling and element-wise addition for simplicity. This efficient and recursive design enables the model to encapsulate multi-frequency representations within each block, effectively replicating the receptive field of large-kernel convolutions without the associated parameter inflation and computational overhead. Furthermore, this flexible and modular design philosophy is compatible with diverse operators, and can be seamlessly integrated into existing vision architectures or extended to other modalities.
Algorithm 1 RecConv in a PyTorch-like style
class RecConv(Module):
def __init__(self, ic, ks=5, level=1):
super().__init__()
self.level = level
kwargs = {
’in_channels’: ic,
’out_channels’: ic,
’kernel_size’: ks,
’padding’: ks // 2,
’groups’: ic
# downsample with shared parameters
self.down = Conv(stride=2, **kwargs)
# convolutions on different levels
self.conv = [Conv(**kwargs)] * (level+1)
def forward(self, x):
i, fs = x, []
for _ in range(self.level):
x, s = self.down(x), x.shape[2:]
fs.append((x, s))
x = 0
for c, (f, s) in zip(self.conv, reversed(fs)):
x = interpolate(c(f + x), size=s)
return self.conv[self.level](i + x)
Specifically, as demonstrated in Algorithm 1 and Equation 3, feature maps of level $\ell$ are produced by the previous level $\ell-1$ using a depthwise convolution with a stride of $2$. This process then generates new feature maps at level $\ell+1$ using the same downsampling operation. Higher-level feature maps are processed through depthwise convolutions and then progressively integrated back into previous levels through bilinear interpolation and element-wise addition. This recursive design maintains parameter efficiency as decomposition levels increase by reusing the downsampling layer and employing a parameter-free upsampling operation across all decomposition levels.
$\begin{split}X_{\ell}&=\mathrm{DWConv_{k\times k,\downarrow}}(X_{\ell-1}),\\ X_{\ell}&=X_{\ell}+\mathrm{Interp_{\uparrow}}\left(\mathrm{DWConv_{k\times k}^{(\ell+1)}}(X_{\ell+1})\right),\end{split}$ (3)
where $X_{\ell-1}\in\mathbb{R}^{B\times C\times 2h\times 2w}$, $X_{\ell}\in\mathbb{R}^{B\times C\times h\times w}$, and $X_{\ell+1}\in\mathbb{R}^{B\times C\times\frac{h}{2}\times\frac{w}{2}}$ represent feature maps at levels $\ell-1$, $\ell$, and $\ell+1$, respectively. $\mathrm{DWConv_{k\times k,\downarrow}}$ indicates the shared-weight depthwise convolution with a kernel size of $k\times k$ and a stride of $2$, while $\mathrm{DWConv_{k\times k}^{(\ell+1)}}$ denotes the depthwise convolution with a kernel size of $k\times k$ at level $\ell+1$. $\mathrm{Interp_{\uparrow}}$ stands for the bilinear interpolation upsampling operation. Features at different scales are progressively fused back into the original feature map, effectively replicating multi-frequency representations with large-kernel receptive fields.
Table 2: Complexity of different types of convolution. The measurement is streamlined by focusing on convolution operations, with the assumption of consistent input and output channels, and excluding the bias term. $k$, $C$, $H$ and $W$ denote the base kernel size, channel number, image height and width, respectively. And $\ell$ represents the number of decomposition levels that determine the effective kernel size, which is $k\times 2^{\ell}$.
Convolution Parameters FLOPs
Standard $\color[rgb]{0.25,0.5,0.5}{k^{2}C}$×$\times(4^{\ell}C)$ $\color[rgb]{0.25,0.5,0.5}{k^{2}CHW}$×$\times(4^{\ell}C)$
Depthwise $\color[rgb]{0.25,0.5,0.5}{k^{2}C}$×$\times(4^{\ell})$ $\color[rgb]{0.25,0.5,0.5}{k^{2}CHW}$×$\times(4^{\ell})$
RecConv $\color[rgb]{0.25,0.5,0.5}{k^{2}C}$×$\times(\ell+2)$ $\color[rgb]{0.25,0.5,0.5}{k^{2}CHW}$×$\times(5/3)$
Table 2 illustrates that to achieve a larger effective kernel size of $k\times 2^{\ell}$, both standard and depthwise convolutions require exponential growth in parameters and FLOPs with respect to the decomposition level $\ell$. In contrast, RecConv exhibits only linear growth in parameters while maintaining constant FLOPs. Specifically, RecConv consists of a shared-weight depthwise convolutional downsampling module and $\ell+1$ depthwise convolution layers, which primarily contribute to the parameter count and FLOPs. The parameter count remains consistent across the downsampling layer and depthwise convolutions at each level, growing by a factor of $\ell+2$ compared to the base kernel. FLOPs can be regarded as a series of geometric progression, decreasing exponentially as the levels increase, as demonstrated in Equation 4. Consequently, RecConv achieves linear parameter growth and nearly constant FLOPs proportional to the decomposition level, making it an efficient solution for attaining large receptive fields in resource-constrained scenarios.
$1+2\sum_{n=1}^{\ell}\frac{1}{4^{n}}<\sum_{n=0}^{\infty}2\cdot\left(\frac{1}{4}\right)^{n}-1=\frac{5}{3}$ (4)
In addition, our RecConv embodies a versatile design philosophy that accommodates different operators and dimensions. For instance, the shared downsampling layer can be substituted with a sequence of channel-wise chunking operations followed by depthwise convolutions; depthwise convolutions at each level can be replaced by efficient attention mechanisms; bilinear upsampling operations can be converted to transposed convolutions; and element-wise addition can be reconfigured as channel-wise concatenation.
4 Experiments
We demonstrate the versatility and efficacy of RecNeXt across multiple vision tasks: classification on ImageNet-1K [8], object detection and instance segmentation on MS-COCO 2017 [35], and semantic segmentation on ADE20K [74]. Following prior works [32, 34, 61, 43, 64, 72], we export the model using Core ML Tools [1] and assess its latency on an iPhone 13 (iOS 18) using the Xcode performance tool. Additionally, we measure throughput on an Nvidia RTX3090 GPU with the maximum power-of-two batch size that can be accommodated in memory.
4.1 Image Classification
Implementation details.
We perform image classification on ImageNet-1K using a standard image resolution of 224$\times$224 for both training and evaluation. This dataset consists of 1.3M training, 50k validation, and 100k test images across 1000 categories. All models are trained from scratch for 300 epochs following the training protocol in [60, 34, 64, 72]. For fair comparisons, we employ the RegNetY-16GF [47] model (top-1 accuracy 82.9%) as the teacher model for hard knowledge distillation. Latency is measured on an iPhone 13 with models compiled via Core ML Tools at a batch size of 1. Performance with and without distillation is reported in Table 1 and 3, respectively.
Table 3: Results without distillation on ImageNet-1K, where“$\dagger$” denotes the evaluation image size is 256.
Model Latency Params Top-1 Auxiliary
(ms) (M) (%) NAS SRP
EfficientFormerV2-S0 [34] 1.4 3.5 73.7
FastViT-T8$\dagger$ [60] 1.3 3.6 75.6
MobileOne-S1 [61] 1.2 4.8 75.9
RepViT-M0.9 [64] 1.3 5.1 77.4
RepNeXt-M1 [72] 1.3 4.8 77.5
RepViT-M1.0 [64] 1.4 6.8 78.6
RepNeXt-M2 [72] 1.4 6.5 78.9
RecNeXt-M1 1.4 5.2 78.0
RecNeXt-M2 1.5 6.8 79.2
MobileOne-S2 [61] 1.5 7.8 77.4
EfficientFormerV2-S1 [34] 1.6 6.1 77.9
MobileOne-S3 [61] 1.7 10.1 78.1
FastViT-T12$\dagger$ [60] 1.8 6.8 79.1
RepViT-M1.1 [64] 1.5 8.2 79.4
RepNeXt-M3 [72] 1.5 7.8 79.4
RecNeXt-M3 1.6 8.2 79.6
MobileOne-S4 [61] 2.2 14.8 79.4
FastViT-S12$\dagger$ [60] 2.0 8.8 79.8
PoolFormer-S24 [68] 3.2 21.0 80.3
EfficientFormerV2-S2 [34] 2.3 12.6 80.4
FastViT-SA12$\dagger$ [60] 2.3 10.9 80.6
RepViT-M1.5 [64] 1.9 14.0 81.2
RepNeXt-M4 [72] 1.9 13.3 81.2
RecNeXt-M4 2.4 14.1 81.1
PoolFormer-S36 [68] 4.3 31.0 81.4
RepViT-M2.3 [64] 2.8 22.9 82.5
FastViT-SA24$\dagger$ [60] 3.4 20.6 82.6
RepNeXt-M5 [72] 2.8 21.7 82.4
RecNeXt-M5 3.4 22.9 81.6
Table 4: Object detection and instance segmentation were evaluated using Mask R-CNN on MS-COCO 2017. Semantic segmentation results were obtained on ADE20K. Backbone latencies were measured on an iPhone 13 with 512$\times$512 image crops using Core ML Tools. Models marked with $\dagger$” were initialized with weights pretrained for 450 epochs on ImageNet-1K.
Backbone Latency $\downarrow$ (ms) Object Detection Instance Segmentation Semantic
[…]
ResNet18 [18] 3.8 34.0 54.0 36.7 31.2 51.0 32.7 32.9
PoolFormer-S12 [68] 6.2 37.3 59.0 40.1 34.6 55.8 36.9 37.2
FastViT-SA12 [60] 7.6 38.9 60.5 42.2 35.9 57.6 38.1 38.0
RepViT-M1.1 [64] 4.1 39.8 61.9 43.5 37.2 58.8 40.1 40.6
RepNeXt-M3 [72] 4.3 40.8 62.4 44.7 37.8 59.5 40.6 40.6
RecNeXt-M3 5.2 41.7 63.4 45.4 38.6 60.5 41.4 41.0
PoolFormer-S24 [68] 10.5 40.1 62.2 43.4 37.0 59.1 39.6 40.3
PVT-Small [66] 26.2 40.4 62.9 43.8 37.8 60.1 40.3 39.8
SwiftFormer-L1 [51] 6.2 41.2 63.2 44.8 38.1 60.2 40.7 41.4
RepViT-M1.5 [64] 5.7 41.6 63.2 45.3 38.6 60.5 41.5 43.6
FastViT-SA24 [60] 12.2 42.0 63.5 45.8 38.0 60.5 40.5 41.0
RepNeXt-M4 [72] 6.1 42.9 64.4 47.2 39.1 61.7 41.7 43.3
RecNeXt-M4 7.6 43.5 64.9 47.7 39.7 62.1 42.4 43.6
SwiftFormer-L3 [51] 10.4 42.7 64.4 46.7 39.1 61.7 41.8 43.9
FastViT-SA36 [60] 16.2 43.8 65.1 47.9 39.4 62.0 42.3 42.9
RepViT-M2.3$\dagger$ [64] 9.0 44.6 66.1 48.8 40.8 63.6 43.9 46.1
RepNeXt-M5 [72] 9.3 44.7 66.0 49.2 40.7 63.5 43.6 45.0
RecNeXt-M5 12.4 44.6 66.3 49.0 40.6 63.5 43.5 46.0
Results with knowledge distillation.
As demonstrated in Table 1, RecNeXt consistently delivers superior performance and competitive latency across various model sizes without requiring additional training auxiliaries. Notably, smaller RecNeXt models outperform their RepNeXt counterparts by 0.2% in top-1 accuracy while maintaining comparable latency and FLOPs. RecNeXt-M5 matches the top-1 accuracy (83.3%) of EfficientFormerV2-L with fewer parameters and lower latency. Without leveraging structural reparameterization (SRP) or neural architecture search (NAS), the RecNeXt series consistently surpasses SwiftFormer in top-1 accuracy with similar computational complexity. However, RecNeXt encounters challenges with throughput due to the extensive employment of bilinear upsample operations, which are less friendly to GPU acceleration. This deficiency can be alleviated by leveraging transposed depthwise convolutions at the cost of increased computational complexity and parameter count.
Results without knowledge distillation.
Table 3 illustrates that RecNeXt maintains strong performance across various model scales. RecNeXt-M1 achieves a top-1 accuracy of 78.0% with a latency of 1.4 ms and 5.2M parameters, outperforming prior lightweight models such as RepViT-M0.9 (77.4%) and RepNeXt-M1 (77.5%) at comparable latencies. RecNeXt-M2 further improves to 79.2% accuracy with 6.8M parameters and 1.5 ms latency, surpassing RepViT-M1.0 (78.6%) and RepNeXt-M2 (78.9%) under similar resource constraints. For higher-capacity models, RecNeXt-M3 achieves a top-1 accuracy of 79.6%, exceeding other leading models without relying on structural reparameterization (SRP) and neural architecture search (NAS). However, larger models like RecNeXt-M4 and RecNeXt-M5 exhibit performance saturation alongside increased latency, with RecNeXt-M5 plateaus at 81.6% top-1 accuracy but with a latency of 3.4ms (0.6ms higher than RepNext-M5). Although RecNeXt-M5 experiences performance degradation, it still outperforms PoolFormer-S36 (81.4%) under the same training protocol, with fewer parameters and lower latency. Additionally, incorporating SRP or advanced operators could potentially mitigate this limitation.
4.2 Downstream Tasks
Object Detection and Instance Segmentation.
We evaluate RecNeXt’s transfer ablility on object detection and instance segmentation tasks. Following [34], we integrate RecNeXt into the Mask-RCNN framework [19] and conduct experiments on the MS-COCO 2017 dataset [35], presenting metrics of APbox and APmask. As shown in Table 4, RecNeXt demonstrates superior performance in terms of APbox and APmask while maintaining competitive latency. For instance, RecNeXt-M3 achieves a notable APbox of 41.7 and APmask of 38.6 while maintaining a latency of 5.2ms, offering an optimal trade-off between speed and precision. RecNeXt-M4 surpasses FastViT-SA24 by 1.5 APbox and 1.7 APmask with 4ms latency improvement, and increases the APbox and APmask by 0.6 compared to RepNext-M4 with only 0.5ms latency increase. However, RecNeXt-M5 encounters an performance saturation at 44.6 APbox and 40.6 APmask, albeit with a higher latency of 12.4ms compared to 9.3ms of RepNext-M5.
[…]
Image Size Levels Kernel Size Receptive Field Params MACs Latency Throughput Top1 Upsample
(M) (G) (ms) (im/s) (%) Operation
$224\times 224$ $3\times 3$ [3, 3, 3, 3] 4.80 0.82 1.24 4479 74.64 Bilinear Interpolate
$7\times 7$ [7, 7, 7, 7] 4.96 0.87 1.27 3529 75.52
$11\times 1$ 11 [11, 11, 11, 11] 5.26 0.96 1.31 2591 75.41
[…]
$5\times 5$ [80, 40, 20, 10] 5.16 0.87 1.36 384 76.03
$7\times 7$ [112, 56, 28, 14] 5.55 0.92 1.43 374 75.67
$384\times 384$ $3\times 3$ [3, 3, 3, 3] 4.80 2.40 2.07 1553 76.35
$7\times 7$ [7, 7, 7, 7] 4.96 2.55 2.20 1224 77.42
$11\times 1$ 11 [11, 11, 11, 11] 5.26 2.82 2.27 880 78.00
[…]
$5\times 5$ [80, 40, 20, 10] 5.16 2.54 2.79 311 78.11
$7\times 7$ [112, 56, 28, 14] 5.55 2.69 2.84 295 78.03
$3\times 3$ [48, 24, 12, 6] 4.97 2.44 2.40 993 77.26 DWConv Transpose
$5\times 5$ [80, 40, 20, 10] 5.31 2.57 2.48 784 77.87
$7\times 7$ [112, 56, 28, 14] 5.81 2.75 2.63 576 78.24
[…]
Semantic Segmentation.
We perform semantic segmentation experiments on the ADE20K dataset [74], which consists of approximately 20K training images and 2K validation images across 150 categories. Intersection-over-Union (mIoU) is used as the evaluation metric. We followed the training protocol from previous works [32, 34] using the Semantic FPN framework [26]. As illustrated in Table 4, RecNeXt demonstrates favorable mIoU-latency trade-offs across various model sizes, with RecNeXt-M3 achieving the highest mIoU of 41.0% among other models of similar scale. RecNeXt-M4 and RecNeXt-M5 also achieve competitive mIoU and latency compared to the state-of-the-art models. These findings emphasize the critical role of efficient backbones with large receptive fields in advancing semantic segmentation, especially in computationally constrained environments.
4.3 Ablation Studies
We systematically investigate the impact of convolutional kernel sizes, recursive decomposition, upsampling operations, and RecConv variants on model performance across different image resolutions. Experiments were conducted over 120 epochs on ImageNet-1K [8] using RecNeXt-M1 as the baseline, at resolutions of $224\times 224$ and $384\times 384$.
Kernel size. Increasing convolution kernel sizes significantly enhance model performance, especially for higher-resolution images. As shown in Table 5, expanding the kernel from $3\times 3$ to $7\times 7$ for $224\times 224$ images boosts top-1 accuracy from 74.64% to 75.52% (+0.88%). However, this increase comes with a rise in parameter count from 4.80M to 4.96M and a slight increase in latency from 1.24ms to 1.27ms. Further enlarging the kernel to $11\times 1$ 11 slightly reduces accuracy to 75.41%, while increasing parameters to 5.26M and rising latency to 1.31ms. For $384\times 384$ images, kernel enlargement delivers consistent accuracy improvements, achieving a +1.65% enhancement (76.35% → 78.00%) as the kernel grows from $3\times 3$ to $11\times 1$ 11, with a modest latency increase (+0.20ms). These findings reveal a trade-off between the computational efficiency of smaller convolution kernels and the enhanced spatial feature capture of larger convolution kernels.
Recursive decomposition. The integration of recursive decomposition expands the model’s receptive field (7→80) and improves multi-scale feature aggregation, leading to accuracy gains across resolutions. For $224\times 224$ images, decomposition with $5\times 5$ kernels achieves a top-1 accuracy of 76.03%, outperforming single-level configurations with larger kernel sizes of $7\times 7$ (75.52%) and $11\times 1$ 11 (75.41%). The benefits are even more pronounced for larger image resolutions ($384\times 384$), where the peak top-1 accuracy of 78.11% is achieved compared to 77.42%, without substantial changes in parameter count or FLOPs. However, the method increases latency (2.2ms $\rightarrow$ 2.8ms) and significantly reduces throughput (1200 img/s $\rightarrow$ 300 img/s). This partially due to the extensive employment of bilinear upsampling operation, which can be replaced by a transposed depthwise convolution for improved performance, albeit at the cost of increased computational complexity. In summary, the progressive decomposition allows the network to effectively aggregate multi-frequency representations while balancing computational complexity by distributing operations across different spatial scales.
Upsampling operation. Table 5 demonstrates the effectiveness of substituting bilinear upsampling with transposed depthwise convolution for $384\times 384$ images, While this replacement slightly increases computational complexity and parameter count, it potentially enhances feature fidelity during upsampling, resulting in a modest increase in top-1 accuracy from 78.03% to 78.24% with $7\times 7$ kernels. Additionally, it reduces latency from 2.84ms to 2.63ms and boosts throughput from 300img/s to 580img/s. However, smaller kernel sizes (e.g., $5\times 5$ and $3\times 3$) lead to performance degradation, with accuracy dropping from 78.11% to 77.87% and from 77.72% to 77.26%, respectively.
RecConv Variants. We present two RecConv variants (Figure 3 and 4) incorporating recursive decomposition along both spatial and channel dimensions. These variants improve computational efficiency but at the cost of accuracy. As summarized in Table 5, the variant in Figure 3 reduces parameter from 5.81M to 5.76M, FLOPs from 2.75G to 2.70G, and increases throughput from 576img/s to 636img/s, but results in a 0.33% drop in accuracy (from 78.24% to 77.91%). Meanwhile, the latency increases significantly due to the lack of adaptation of the transposed group convolution on the NPU. The variant in Figure 4 further reduces parameters by 0.4M and FLOPs by 0.1G, lowers latency by 0.3ms, and increases throughput by 200 img/s, but causes a more significant accuracy drop of 1.4%. This comparative analysis highlights the trade-off between upsampling strategies and decomposition methods in optimizing performance and computational efficiency. This study emphasizes design simplicity, using parameter-free bilinear upsampling operation and spatial-only decomposition to achieve a balanced solution.
Figure 3: RecConv Variant A. Recursive decomposition is implemented along both spatial and channel dimensions simultaneously. The shared downsample operation is replaced with group convolutions, where the number of groups matches the number of output channels. Bilinear upsample operations are substituted with transposed group convolutions, where the number of groups corresponds to the number of input channels.
Figure 4: RecConv Variant B. The recursive decomposition operation is jointly performed in both spatial and channel dimensions. The shared downsample operation is replaced with channel-wise split operations, followed by depthwise convolutions. Bilinear upsample operations are converted to transposed depthwise convolutions, and element-wise addition is substituted with channel-wise concatenation.
5 Conclusions
This paper introduces a straightforward and versatile recursive decomposition strategy that leverages small-kernel convolutions to construct multi-frequency representations, establishing a linear relationship between parameter growth and decomposition levels. This innovation guarantees that the computational complexity at each decomposition level follows a geometric progression, diminishing exponentially with increasing depth. This recursive design is compatible with various operations and can be extended to other modalities, though this study focuses on vision tasks using basic functions. Leveraging this approach, we construct RecConv as a plug-and-play module that can be seamlessly integrated into existing computer vision architectures. To the best of our knowledge, this is the first effective convolution kernel up to $80\times 80$ for resource-constrained vision tasks. Building on RecConv, we introduce RecNeXt, an efficient large-kernel vision backbone optimized towards resource-constrained scenarios. Extensive experiments across multiple vision benchmarks demonstrate RecNeXt’s superiority over current leading approaches without relying on structural reparameterization or neural architecture search.
Limitations: While RecNeXt delivers competitive performance, its scalability lags behind the state-of-the-art models without knowledge distillation, where larger variants experience performance saturation and increased latency. Additionally, RecNeXt exhibits the lowest throughput among models of comparable parameter size due to the computational overhead of bilinear interpolation, which can be mitigated by employing transposed convolution. The recursive decomposition may introduce numerical instability during mixed precision training, which can be alleviated by using fixed-point or BFloat16 arithmetic. Compatibility issues with bilinear interpolation and transposed convolution on certain iOS versions may also result in performance degradation.